OpenStudy (anonymous):

will give metal for answer and explination For the function f(x) = 2 – 3x, find f(0). A. 2 B. –1 C. 5 D. –3

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

replace the x in the formula with 0 then simplify

OpenStudy (anonymous):

B

OpenStudy (anonymous):

show me how you got it

OpenStudy (anonymous):

2/3=-1

OpenStudy (anonymous):

The equation is 2 - 3x. Where did 2/3 come from?

OpenStudy (anonymous):



OpenStudy (anonymous):

i must be thinking of something alse

OpenStudy (anonymous):

else

OpenStudy (anonymous):

so then 2-3=1*0=-1 that's what i'm coming up with

OpenStudy (anonymous):

order of operations. Parentheses Exponents Multiplication and Division Addition and Subtraction